OpenStudy (cutiekawdfi):

factor perfect square trinomial x^2-14x+49 factor difference of two squares 36x^2-16

OpenStudy (alexandervonhumboldt2):

(a-b)^2=a^2-2ab+b^2 look: x^2-2*7*1+7^2=(x-7)^2

OpenStudy (lochana):

hint for second one : \[(x^2-a^2) = (x-a)(x+a)\]

OpenStudy (lochana):

@cutiekawdfi have you found solutions?

OpenStudy (lochana):

\[(36x^2 - 16) = ((6x)^2 - 4^2)\]